An electroplating unit plates 3.0 g of silver on a brass plate in 3.0 minutes. Find the current used by the unit. The electrochemical equivalent of silver is 1.2 × 10^-6 kg/C.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. t = 3 min = 180 sec w = 2 g E.C.E = 1.12 × 10^–6 kg/c ⇒ 3 × 10–3 = 1.12 × 10^–6 × i × 180 ⇒ i = 3 * 10^-3/1.12 *10^-6 *180 = 1/6.72 *10^2 = 15 Amp.